If a:b:c is 8:14:22 and b:c:d is 21:33:44 so what is a:b:c:d ?
Answer by Theo(13342)

the ratios are:

first set:
a = 8
b = 14
c = 22
d = ????

second set:

a = ????
b = 21
c = 33
d = 44

if these ratios are in proportion, then:

a2/a1 should be equal to b2/b1 should be equal to c2/c1 should be equal to d2/d1

you have 2 ratios to work from:

b2/b1 = 21/14 = 1.5

c2/c2 = 33/22 = 1.5

looks like the ratios are the same, so the ratios from the first set are probably in proportion to the ratios in the second set.

in order for that to hold true, then:

a2/a1 would also have to be equal to 1.5

d2/d1 would also have to be equal to 1.5

a2/a1 is equal to a2/8

for a2/8 to be equal to 1.5, then a2 would have to be equal to 1.5 * 8 = 12.

d2/d1 is equal to 44/d1

for d2/d1 to be equal to 1.5, then d1 would have to be equal to 44 / 1.5 = 29 and 1/3.

the complete ratios for both sets would be:

first set:

a:b:c:d = 8:14:22:29 and 1/3

second set:

a:b:c:d = 12:21:33:44

the ratios a2/a1, b2/b1, c2/c1, d2/d1 would all have to be the same, as they are.

the following ratios would also have to be equal:

b1/a1 = b2/a2
c1/b1 = c2/b2
d1/c1 = d2/c2

as they are.

just in case, you didn't catch it, b1 is b from set 1 and b2 is b from set 2, etc.